### v4

# TextCenGen: Attention-Guided Text-Centric Background Adaptation for Text-to-Image Generation[Uncaptioned image]

Abstract
Recent advancements in Text-to-Image (T2I) generation have witnessed a shift from adapting text to fixed backgrounds to creating images around text. Traditional approaches are often limited to generate layouts within static images for effective text placement. Our proposed approach, TextCenGen, introduces a training-free and plug-and-play dynamic background adaptation of the blank region for text-friendly image generation. Our method employs force-directed attention guidance in T2I models to generate images that strategically reserve whitespace for any pre-defined text areas. Observing how cross-attention maps affect object placement, we detect and repel conflicting objects using a force-directed graph approach, combined with a spatial excluding cross-attention constraint for smooth attention in whitespace areas. To evaluate our method, we propose hybrid prompt datasets as benchmarks with . Experiments indicate that TextCenGen outperforms existing methods by producing more harmonious compositions, demonstrating that it effectively enhances T2I model outcomes by catering to different text positions.
Figure 1: TextCenGen is a training-free method designed to generate text-friendly images. By using a simple text prompt and a planned blank region as inputs, TextCenGen creates images that satisfy the prompt and provide sufficient blank space in the target region. For example, the text-friendly T2I approach helps users customize their favored text-friendly wallpapers for mobile devices with any T2I model, avoiding visual confusion caused by main objects overlapping with UI components.
Introduction
111*These authors contributed equally to this work222†Corresponding author: chli@cs.ecnu.edu.cn
In graphic design, achieving an effective synergy between text and imagery is essential for clear communication. The choice of background can influence text visibility and comprehension. A common design objective is to place text within an image in a way that is both visually pleasing and communicatively clear. A preferred strategy is to position the text in the golden ratio, which is believed to be aesthetically optimal. However, designers often grapple with the issue of backgrounds that compete with or obscure the text, as illustrated in Figure 1[Uncaptioned image], where unsuitable backgrounds detract from the text’s readability and aesthetic appeal, regardless of any adjustments to text color or size. We aim to facilitate the creation of text-friendly images (see Figure 1[Uncaptioned image]), ideal for text placement and meeting the growing demand due to increasing use of Text-to-Image (T2I) models for background graphics.
Traditional approaches to graphic design, especially in poster creation, have largely focused on arranging layouts in static natural background images, elements, and text (Guo et al. 2021; Cao, Chan, and Lau 2012; O’Donovan, Agarwala, and Hertzmann 2014; Li, Zhang, and Wang 2022). However, producing text-friendly images remains a significant challenge due to the complexity of background elements. Our insight, derived from Figure 1, reveals that a clear separation between the main objects and the text areas is essential. Recent advancements in diffusion-related research have shown that it is possible to manipulate the primary objects within cross-attention maps, making the adaptation of background images to accommodate text a feasible endeavor (Hertz et al. 2022; Epstein et al. 2023).
In response, we introduce TextCenGen[Uncaptioned image], as illustrated in Figure 1, a novel method that employs cross-attention maps and force-directed graphs for effective object placement and whitespace optimization. We also implement a spatial excluding cross-attention constraint to ensure smooth attention in areas designated for text. To establish a new benchmark for this innovative task, we constructed a diverse dataset gathered from three unique sources, along with three distinct evaluation metrics, to comprehensively assess the performance. The contributions of our paper are three-fold:
  • We propose a new task of text-freindly T2I generation, which creates images that satisfy both the prompt and reserve space for pre-defined text placements. The task consists of a benchmark including a specialized dataset and tailored evaluation metrics.
  • We introduce TextCenGen, a plug-and-play, training-free framework. This approach allows for dynamic placement of text at user-specified location within an image, with the image being generated or adapted to accommodate the text.
  • Our prominent module, the force-directed cross-attention guidance, is an effective addition to the denoising process in T2I models. This module strategically directs cross-attention map within the image generation process, ensuring harmonious layout of text and imagery.
Related Work
Text Layout of Natural Images has evolved significantly, transitioning from traditional layout designs to more advanced methods influenced by deep learning. Initially, poster design focused on creating layouts with given background images, elements, and text (Guo et al. 2021; Cao, Chan, and Lau 2012; O’Donovan, Agarwala, and Hertzmann 2014; Li, Zhang, and Wang 2022). The integration of deep learning in text layout of natural image generation has led to various models such as GAN (Goodfellow et al. 2014; Zheng et al. 2019; Li et al. 2019; Zhou et al. 2022), VAE (Jyothi et al. 2019), transformers (Vaswani et al. 2017; Inoue et al. 2023b; Wang et al. 2023) and diffusion models (Ho, Jain, and Abbeel 2020; Hui et al. 2023; Inoue et al. 2023a; Chai, Zhuang, and Yan 2023; Li et al. 2023a). These models have been instrumental in learning layout patterns from large datasets. Subsequent research explored methods to retrieve matching background images based on text and image description (Jin et al. 2022). Since the development of image editing method based on diffusion model, scene text generation methods such as TextDiffuser (Chen et al. 2023b, a) and DiffText (Zhang et al. 2023) have addressed the challenges of generating clear text with diffusion models. However, these methods often rely on the presence of a “sign” or similar element within the prompt (e.g., a T-shirt) to place text. They do not explicitly tackle the problem of creating text-friendly images where the background itself is crafted to adapt pre-defined text regions. Our approach extends these capabilities by allowing the primary objects in generated images to yield space to text regions, resulting in more harmonious and aesthetically pleasing compositions.
Text-to-Image Generation has advanced with diffusion models (Ho, Jain, and Abbeel 2020), producing realistic images and videos that align closely with text prompts (Rombach et al. 2022; Singer et al. 2023; Chen et al. 2023c). Innovations in this field include GLIDE (Nichol et al. 2022), which integrates text conditions into the diffusion process, Dall-E 2 (Ramesh et al. 2022) with its diffusion prior module for high-resolution images, and Imagen (Saharia et al. 2022), which uses a large T5 language model to enhance semantic representation. Stable diffusion (Rombach et al. 2022) projects images into latent space for diffusion processing. Beyond text conditions, the manipulation of diffusion models through image-level conditions has been explored. Methods such as image inpainting (Balaji et al. 2022) aim to generate coherent parts of an image, while SDG (Liu et al. 2023) introduces semantic inputs to guide unconditional DDPM sampling. In addition, techniques such as (Meng et al. 2021) use images as editing conditions in denoising processes. Scene text generation methods such as TextDiffuser (Chen et al. 2023b, a) and GlyphDraw (Ma et al. 2023) have also emerged, utilizing textual layouts or masks to guide text generation in images. These advancements represent the growing versatility and potential of T2I models in diverse applications.
Attention Guided Image Editing has emerged as a foundamental solution to the challenge of translating human preferences and intentions into visual content through text descriptions. These approaches, as seen in works like (Li et al. 2023b; Avrahami et al. 2023; Zhang, Rao, and Agrawala 2023; Zhao et al. 2023), involve learning auxiliary modules on paired data. However, a significant limitation of these training-based methods is the substantial cost and effort required for repeated training for different control signals, model architectures, and checkpoints. In response to these challenges, training-free techniques have emerged, using the inherent weights of attention and the characteristics of pre-trained models to control the characteristics of objects such as size, shape, appearance, and location (Hertz et al. 2022; Epstein et al. 2023; Patashnik et al. 2023; Xie et al. 2023). These methods typically utilize basic conditions, such as bounding boxes, for precise control over object positioning and scene composition (Mo et al. 2023). Desigen (Weng et al. 2024) discovered relationship between attention and saliency and introduce attention reduction to weaken the attention whin layout boxes. Our approach takes this further by applying force-directed graph techniques to cross-attention map edits, allowing for more automated and precise object transformations in T2I editing.
Method
Figure 2: In our approach, the model receives a blank region ($R$) denoted as red-dotted area, and a text prompt as its inputs. The prompt is then used concurrently in a Text-to-Image (T2I) model to generate both an original image and a result image. During each step of the diffusion model’s denoising process, the cross-attention map from the U-Net associated with the original image is used to direct the denoising of the result image in the form of a loss function. Throughout this procedure, a conflict detector identifies objects that could potentially conflict with $R$. To mitigate such conflicts, a force-directed graph method is applied to spatially repel these objects, ensuring that the area reserved for the text prompt remains unoccupied. To further enhance the smoothness of the attention mechanism, a spatial excluding cross-attention constraint is integrated into the cross-attention map.
Given an input text region $R$ and a prompt $T_{prompt}$, our framework aims to generate a text-friendly image $I_{res}$. Specifically, the output image has (1) reduced overlap between the primary object and $R$, (2) sufficient smoothness and minimal color variation in $R$, and (3) the image still fits the prompt. Our framework is shown in Figure 2.
As cross-attention map $A_{k}$ serves as a medium to locate and edit objects within generated images (Epstein et al. 2023), we focus on the $k^{th}$ token of $T_{prompt}$. In response to our first concern, we analyze the denoising process of an unguided image ($I_{ori}$) from a given prompt, and establish the subset of tokens ($O$) that refer to objects that need modification. We design a conflict detector that determines object conflicts based on the average attention intensity in the overlapping regions between $A_{k}$ and $R$. For every token $k\in O$, we introduce Force-Directed Cross-Attention Guidance for moving objects. In this scheme, objects are treated as centroid vertex within a graph, with a sequence of forces being applied (see Figure 3) to adjust object positions.
For the second consideration, inspired by recent technologies that limit the range of the attention map (Zhang et al. 2023), we propose a spatial excluding cross-attention constraint to prevent an extensive attention density from encroaching on $R$.
Addressing the third concern, our denoising process incorporates a loss function with additional regularization terms to safeguard the shapes and positions of other objects. Sometimes excessive repulsive force can occasionally displace essential objects from the image. To prevent objects from being dislocated outside limits while retaining their reasonable shapes, we also introduce the notions of Margin Force $F_{m}()$ and Warping Force $F_{w}()$.
Force-Directed Cross-Attention Guidance
Figure 3: Illustration of four set relationships and their associated forces. The Repulsive Force separates object and text centroids during intersections (a1) and enclosures (a2). The Margin Force (b) and Warping Force (c) prevent boundary overstepping. Text within object regions (a4) requires cooperation between force and attention constraint.
Cross-Attention and Centroid of Object.
To seamlessly integrate the concept of force-directed graphs into the loss guidance of the denoising process in latent diffusion models, we delve into the extraction and manipulation of attention maps and activations. We use softmax normalized attention matrices $\mathcal{A}_{i,t}\in\mathbb{R}^{H_{i}\times W_{i}\times K}$ of layer $i$ extracted from the standard denoising forward step $\epsilon_{\theta_{i}}(z_{i};t,y)$. This enables us to manipulate the control over objects referred to in the text conditioning $y$ at distinct indices $k$, by adjusting the related attention channel(s) $\mathcal{A}_{i,t,\ldots,k}\in\mathbb{R}^{H_{i}\times W_{i}\times|k|}$. The centroid of the attention map is a two-dimensional vector, defined by the equation:
$\text{centroid}\left(k\right)=\frac{1}{\sum_{h,w}\mathcal{A}_{h,w,k}}\begin{bmatrix}\sum_{h,w}h\mathcal{A}_{h,w,k}\\ \sum_{h,w}w\mathcal{A}_{h,w,k}\end{bmatrix}.$ (1)
To ensure the efficacy of our centroid-based object representation method, we initially assume that all objects are convex sets, adhering to the mathematical definition that for every pair of points within the object, the line segment connecting them lies entirely within the object. This assumption allows us to treat the extracted $\text{centroid}\left(k\right)$ as vertices $v_{k}$ in a graph, which are then subjected to force-directed attention guidance. Indeed, it is important to clarify that while each token $k$ associated with $\mathcal{A}_{k}$ appears as a single entity in Figure 2, it actually represents an average of $\mathcal{A}_{k}^{l}$ across all layers $l$ in the U-Net architecture. Practically, our method is applied individually to each layer, ensuring a nuanced and layer-specific approach to force-directed attention guidance.
Layer-wise Conflict Multi-Target Detector.
To effectively manage conflicts between text and objects in our images, we have developed a layer-wise conflict multi-target detector, denoted as $D()$. This detector is crucial for identifying tokens $k$ within each layer $l$ of the U-Net that correspond to objects that require modifications in relation to text regions. The detector function $D(k,R,A_{k}^{l})$ operates as follows:
$D(k,R,A_{k}^{l})=\begin{cases}1,&\text{if mean}(\mathcal{A}_{h,w,k}^{l}\cap R)>\theta\\ 0,&\text{otherwise}\end{cases}$ (2)
where $\mathcal{A}_{h,w,k}^{l}$ is the attention map for token $k$ at layer $l$, and $R$ is the region designated for text. The function returns a value of 1 when the mean attention within the overlap between $\mathcal{A}_{h,w,k}^{l}$ and $R$ exceeds a predefined threshold $\theta$, indicating a conflict that requires our guidance function. Identifying and adjusting the bounding boxes is visualized in Figure 2.
Repulsive Force.
The fundamental repulsive force $F_{rep}(v_{i},v_{j})=\frac{-\xi^{2}}{||pos(v_{i})-pos(v_{j})||}$, ensures that each element is placed separately. $\xi$ denotes the general strength of the force, while $p(v_{i})$ and $p(t/o)$ indicate the positions of the vertex and the target object, respectively. For scenarios that encompass multiple targets, our framework adopts a cumulative force approach to balance attention across these elements. This is quantified by the formula:
$F_{multi-tar}(v_{i})=\sum_{j=1}^{n}\omega_{j}\cdot\frac{-\xi^{2}}{||p(v_{i})-p(tar_{j})||},$ (3)
where $\omega_{j}$ are coefficients for balancing attention across targets. To regulate the impact of these forces and avoid excessive dominance by any single target, we introduce a force balance constant $\alpha$ in the form of $\frac{F_{rep}}{\alpha+F_{rep}}$. $\alpha$ ensures that the forces exerted do not exceed a practical threshold, thereby maintaining visual equilibrium in complex scenes.
Margin Force.
The Margin Force is a critical component of our force-directed graph algorithm, designed to prevent significant vertices from being expelled from visual boundaries. This force, $F_{m}(v)=\frac{-m}{d(v,\text{border})^{2}}$, is activated as a vertex $v$ approaches the edge of the display area, typically a delineated rectangular space. The force is directed inward to ensure that crucial vertices remain within the designated visual region. The constant $m$ modulates the force’s intensity, and $d(v,\text{border})$ represents the distance of the vertex $v$ from the nearest boundary (see Figure 3).
Displacement and Position Update.
To compute the total displacement $\Delta pos(v)$, sum the repulsive and margin forces together: $\Delta pos(v)=F_{rep}(v)+F_{m}(v)$ Subsequently, update the vertex’s position as follows: $pos(v)=pos(v)+\Delta pos(v)$. For the attention map $A_{k}^{l}$, this update is applied as a whole, with excess regions outside the boundaries being discarded and the remaining areas filled with zeros. But this method introduces the risk of the object being moved out of the boundaries and then being discarded, so it is necessary to introduce the following ’warping force’ to prevent this from happening.
Warping Force.
In addressing the dynamics of our force-directed graph algorithm, particularly for the movement of cross-attention maps $\mathcal{A}_{k}^{l}$ in each layer, we employ affine transformations as a pivotal mechanism. This approach facilitates the comprehensive translation and scaling of the entire map, preserving the relative positions within the image domain. Initially, we delineate our visual area or image space as a $H\times W$ two-dimensional array $A$. Within this canvas $A$, we identify a key region, the object $O$, defined by the coordinates $(x,y,a,b)$, where $(x,y)$ marks the upper-left corner and $(a,b)$ the lower-right corner. The movement is calculated based on the sum of repulsive force $F_{\text{rep}}(v)$ and margin force $F_{\text{m}}(v)$, yielding the total displacement $\Delta\text{pos}(v)=F_{\text{rep}}(v)+F_{\text{m}}(v)$. Applying $\Delta\text{pos}(v)$ to both $A$ and $O$, we obtain the transformed canvas $A^{\prime}$ and the object $O^{\prime}$. Subsequently, we shift our coordinate system’s origin to a vertex within $A^{\prime}$ that remains within the canvas boundaries, establishing a new origin $O_{\text{new}}$. This repositioning is crucial when $O^{\prime}$ exceeds the visual boundaries of $A$. In such cases, we scale the moved $\mathcal{A}_{k}^{l}$ to ensure that the bounding box of $O^{\prime}$ fits precisely within the confines of $\mathcal{A}_{k}^{l}$. Scaling factors $S_{x}$ and $S_{y}$, are calculated as $S_{x}=\min\left(1,\frac{H-1}{a^{\prime}}\right)$ and $S_{y}=\min\left(1,\frac{W-1}{b^{\prime}}\right)$, where $(a^{\prime},b^{\prime})$ are the new coordinates of $O^{\prime}$. Finally, the scaled $\mathcal{A}_{k}^{l}$ and $O^{\prime}$ are reverted back to their original coordinate origin (see the warping force in Figure 3). A critical aspect of our approach is the transformation of the reference frame. After displacing $O$ due to $\Delta\text{pos}(v)$, a new reference frame is established, centered at $O_{\text{new}}$. The coordinates of $O$ in this new frame are calculated as $(x_{\text{new}},y_{\text{new}})=(x+\Delta x-O_{\text{new}_{x}},y+\Delta y-O_{\text{new}_{y}})$. The affine transformation, considering this frame shift, is represented as:
$\textbf{T}=\begin{pmatrix}S_{x}&0&\Delta x-O_{\text{new}_{x}}\\ 0&S_{y}&\Delta y-O_{\text{new}_{y}}\\ 0&0&1\end{pmatrix}.$ (4)
This matrix transforms the coordinates of $O$, ensuring that it remains visible within $A$ after transformation. This carefully planned process secures the region $O$ within $A$, even after dynamic changes. It upholds the structure of the cross-attention map and the core principles of our force-directed graph algorithm, balancing key vertices visibility and graph fluidity.
Spatial Excluding Cross-Attention Constraint
Our goal is to maintain a smooth background in the text region denoted as $R$. As illustrated in Figure 2, during each time-step of the forward pass in the diffusion model, we modify the cross-attention maps at every layer. The cross-attention map is represented as $\mathcal{A}_{k}^{l}\in\mathbb{R}^{H\times W\times K}$, where $H\times W$ are the dimensions of $A^{l}$ at different scales, and $K$ signifies the maximum token length at layer $l$. The set $I$ consists of indices of tokens corresponding to areas outside the text in the prompt. We resize $R$ to align with the $HW$ dimensions. Subsequently, a new cross-attention map for each layer $l$ is computed as $\mathcal{A}_{k,\text{new}}^{l}=\{\lambda\times\mathcal{A}_{k}^{l}\odot(1-R)\,|\,\forall k\in O\}$, where $O$ represents the tokens needing editing. This procedure effectively redirects the model’s attention away from $R$, ensuring that the background in this region remains undisturbed and visually smooth. This spatially exclusive approach enhances the clarity and coherence of the generated images, particularly in areas designated for text insertion.
The final loss function of denoising process is defined as:
$\mathcal{L}=||\mathcal{A}_{ori}-\mathcal{A}_{res}||^{2}+\gamma\mathcal{L}_{norm}.$ (5)
In this case, $\mathcal{A}_{ori}$ and $\mathcal{A}_{res}$ stand for the original and altered versions of matrix $A$, correspondingly, while ||.||||.||| | . | | signifies the Euclidean distance among these cross-attention maps. Furthermore, the inclusion of the regularization term $\gamma\mathcal{L}_{norm}$ derived from self-guidance (Epstein et al. 2023), strives to prevent overfitting, where $\gamma$ denotes a weighting parameter.
Figure 4: The results of comparison. Each column showcases six prompts across three datasets, the final column depicting the saliency map of the result image generated from the mushroom prompt. The red-dotted area denotes the planned blank region. Note that some methods fail to follow the orange-highlighted words in the prompt, leading to semantic loss.
Dataset Metrics Stable Diffusion Dall-E 3 AnyText Desigen Ours
P2P Template Saliency IOU $\downarrow$ 29.89 52.64 54.34 30.62 28.37
TV Loss $\downarrow$ 14.11 18.02 22.55 13.70 13.29
VTCM $\uparrow$ 2.95 1.92 1.75 2.92 3.09
DiffuisonDB Saliency IOU $\downarrow$ 30.11 56.82 55.88 30.30 27.80
TV Loss $\downarrow$ 12.30 21.16 20.39 11.99 11.55
VTCM $\uparrow$ 3.19 1.67 1.79 3.20 3.37
Synthetics Prompt Saliency IOU $\downarrow$ 31.42 51.52 53.24 31.57 30.83
TV Loss $\downarrow$ 15.61 17.85 21.46 15.63 15.17
VTCM $\uparrow$ 2.73 1.98 1.79 2.67 2.77
Table 1: Quantitative comparison of our method with other methods (excluding CLIP Score). Bold indicates the best scores.
 Dataset Dall-E 3 AnyText Desigen Ours
P2P Template 2.53 0.32 0.55 0.30
DiffuisonDB 2.04 1.04 0.34 0.31
Synthetics Prompt 2.25 1.02 0.51 0.29
Table 2: Quantitative comparison of CLIP Scores Loss for different methods. Bold indicates the best scores.
Experiments
The evaluation is structured into quantitative and qualitative analyses, alongside an ablation study to understand the contribution of individual components of our model.
Implementation Details.
Experimental Settings.
Our model is built with Diffusers. The pre-trained models are stable-diffusion-v1-5. While generating, the size of the output images is $512\times 512$. We use one A6000 and ten A40 GPUs for evaluation. Detailed parameter settings are provided in the appendix.
Dataset for Evaluation.
Our evaluation contains 27,000 images generated from 2,700 unique prompts, each tested in ten different random region $R$. The dataset combined synthesized prompts generated by ChatGPT, and 700 prompts from the Prompt2Prompt template (Hertz et al. 2022) designed for attention guidance, focusing on specific objects and their spatial relationships. Additionally, we included 1,000 DiffusionDB prompts (Wang et al. 2022), chosen for their real-world complexity. This diverse and comprehensive dataset, spanning synthetic to user-generated prompts, provided a broad test ground to evaluate the efficacy of our TextCenGen method in various T2I scenarios.
Comparison with Existing Methods
We compared TextCenGen with several potential models to evaluate its efficiency. The baseline models included: Native Stable Diffusion (Rombach et al. 2022), Dall-E 3 (Ramesh et al. 2022), AnyText (Tuo et al. 2023) and Desigen (Weng et al. 2024). Dall-E used the prompt ”text-friendly in the position” to specify the region $R$. Similar to AnyText, we chose to randomly generate several masks in a fixed pattern across the map to simulate regions need to be edited. More detail can be found in appendix.
Metrics and Quantitative Analysis.
To evaluate model performance, we used metrics assessing various aspects of the generated images. We proposed the CLIP Score Loss to evaluate the reduction in prompt semantic alignment for the training-free method compared to the original diffusion model. The CLIP Score (Hessel et al. 2021; Huang et al. 2021; Radford et al. 2021) measured semantic fidelity, ensuring images align with textual descriptions. The total variation (TV) loss (Rudin and Osher 1994; Jiang, Li, and Wang 2021) assessed the visual coherence and smoothness of the background in relation to text region $R$, crucial for harmonious compositions. Saliency Map Intersection Over Union (IOU) quantified focus and clarity around text areas. We propose the Visual-Textual Concordance Metric (VTCM) that combines a global metric increasing with value (CLIP Score) and local metrics that benefit from lower values (Saliency IOU and TV Loss) within $R$. The VTCM formula is:
$VTCM=\text{CLIP Score}\times(\frac{1}{\text{Saliency IOU}}+\frac{1}{\text{TV Loss}}).$ (6)
Our quantitative analysis in Table 1 and Table 2 shows that TextCenGen minimizes semantic loss while maintaining smoothness and saliency harmony within region R, even surpassing the latest Dalle-3. Scene text rendering and attention reduction in Desigen often focus on local attention, neglecting global semantics. Especially in Text in Object situation (see Figure 3), this weakens the ability to create whitespace. Our approach moves main objects away, then reduces attention to create space, resulting in more natural and harmonious text layouts with the highest VTCM.
MLLM-as-Judge ELO Ranking.
Following rising trend of multimodal large language model (MLLM) as judge methods (Chen et al. 2024; Wu et al. 2024), we present the MLLM-as-Judge ELO ranking for design appeal across different datasets in Table 3. The results demonstrate that the structured output from the GPT-4o with structured output provides consistent ratings, indicating that our method, along with Anytext and Dall-E, shows advantages in terms of design appeal. Interestingly, Dall-E excels in the LLM dataset that contains a higher proportion of natural landscapes. For detailed information on the prompts and evaluation methodology, please refer to the appendix.
Rank Method DDB P2P LLM
1 TextCenGen 702.21 752.85 122.96
2 Anytext 279.56 329.32 -89.68
3 Dall-E -17.32 -39.05 78.87
4 Desigen -322.63 -291.80 -7.24
5 SD1.5 -629.33 -738.83 -92.40
Table 3: Method ELO design appealing rankings across three datasets.
Qualitative Analysis.
Our qualitative analysis, shown in Figure 4, involved a comparison across different models using prompts and positions within the same quadrant. Dall-E 3, which relies solely on text inputs, exhibited significant variability and could not consistently clear the necessary areas for text. Desigen reduces attention in region $R$, but this method is not always effective without pretrained graphic design-specific model weights, especially when region $R$ is within a main object, such as in the bicycle example. Introducing text to images was tricky, as TV Loss showed, but TextCenGen kept object detail and background quality well. It shows that our force-directed method effectively balances text and visuals in images.
User Study
To understand the importance of the text-friendliness issue and explore users’ subjective perceptions of the T2I Model’s results, we conducted a user study with 111 participants. We used the qualitative results (see Figure 4) to develop a questionnaire. Figure 5 illustrates that the prompt-image alignment and aesthetics of our results were well-received in human perception.
Ablation Study
Table 4 presents the results of our ablation study. As part of our comprehensive analysis, we evaluated the two main contributors: (1) the impact of the force-directed component and (2) the effectiveness of the Spatial Excluding Cross-Attention Constraint.
Impact of Force-Directed Cross-Attention Guidance.
Our model’s force-directed feature is key to gently shifting where objects are placed. Without this, we might just bluntly edit the cross-attention map, which could mess up important parts of the picture. This part of our model helps us make sure we don’t ruin the image structure by harshly removing attention map from areas.
Effects of Spatial Excluding Cross-Attention Constraint.
Despite successful relocation of conflict object-related tokens, spaces left behind may not inherently lead to a well-blended transition. Our experimental results underscore that the integration of the Spatial Excluding Cross-Attention Constraint improve the smoothness of the remaining image sections.
Figure 5: User study of task importance and result evaluation. The left shows user perceptions of task encounter frequency (a1) and importance (a2), rating 5 as the highest. The right side details user rankings (b2) and the average times our results were chosen in a multiple-selection scenario (b1). The y-axis in b2 represents rankings from 1 to 5, demonstrating significant mean differences ($\alpha$=0.05) across three standards for our method.
w/o Both w/o FDG w/o SEC Ours
CLIP Score Loss $\downarrow$ - 2.2 1.71 0.32
TV Loss $\downarrow$ 14.39 12.44 12.76 13.67
Saliency IOU $\downarrow$ 30.32 28.56 28.61 28.34
VTCM $\uparrow$ 2.93 3.03 3.05 3.07
Table 4: Ablation study result. We examine TextCenGen with or without the implementation of Force-Directed Cross-Attention Guidance (FDG) and Spatial Excluding Cross-Attention Constraint (SEC). The CLIP Score Loss of w/o both indicates the use of the original SD-1.5 without our training-free method, resulting in no loss (-).
Conclusion
We present TextCenGen, a novel shift in T2I generation. This model abandons the traditional method of adapting text to pre-defined images for a more flexible, text-focused approach. TextCenGen modifies images to adapt text, employing force-directed cross-attention guidance to arrange whitespace. Furthermore, we have integrated a system to identify and relocate conflicting objects and a spatial exclusion cross-attention constraint for low saliency in whitespace areas.
Our approach has certain limitations. The force-directed cross-attention guidance, assuming convexity and centering on object centroids, may not be suitable for non-convex shapes. It may lead to reduced size or fragmentation of objects. Future work will address these challenges to improve the quality of the output images. Additionally, we observed the generation of unintended objects in blank areas not mentioned in the original prompts.